or the pulley system shown, ignore friction and air resistance. The pulley can be modeled as a disk

  • M = 2 kg
  • m = 0.2 kg
  • R = 0.1 meters

a) What is the magnitude of the acceleration of the block, in m/s2?

b) What is the magnitude of the angular acceleration of the pulley, in rad/s2?

c) What is the magnitude of the torque exerted on the pulley, by the tension in the rope, in N⋅m?

M G O R a T 16 m mg
Answer #1

that given R TY لے m= o.2 kg M=2 R=ollm I= mp for disc 2 kg Now 2 mg mg-T=ma apply forque at point o z=Ish = R XP TR = I=I Ilwhich is the acceleration of of 2 <= 16.666 vod/s31 angular 3 the pully e Torque T-R+T = PT - IL c=1MR =lme) (16-666) = 2(0.1if you have any queries then give me a comment and give me a like thanks.